Please help me find the 50th term in the arithmetic sequence.
5,9,13,17,...... Also, please explain it to me, thank you!

Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39618)

Common difference is 4, and the sequence increases.

Answer by greenestamps(13200)



Use the definition of an arithmetic sequence and do some simple calculations.

The first term is 5.
The second term, 9, is the first term, 5, plus the common difference; so the common difference is 4. Then...

The 3rd term is the 1st term, plus 2 times the common difference: 5+2(4)=5+8=13
The 4th term is the 1st term, plus 3 times the common difference: 5+3(4)=5+12=17
...
The 50th term is the 1st term, plus 49 times the common difference...